OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 2y = 6 x + 3y = −1 −2 2 5 −5

OpenStudy (kamille):

DO you know how to solve the system of eq?

OpenStudy (anonymous):

not really

OpenStudy (kamille):

Well, have you ever heard of any of these methods: systems of eq solving by substitution method and solovning by substraction method?

OpenStudy (anonymous):

uhmm yeaa ive heard of them

OpenStudy (kamille):

So which one do you think you should use here?

OpenStudy (anonymous):

i have no clue. this is my worse subject. maybe substitution method

OpenStudy (kamille):

Well, if you need to find x, which one (y or x) do you need to substitute?

OpenStudy (anonymous):

y ?

OpenStudy (kamille):

yes, you are right. 2x + 2y = 6 y=?

OpenStudy (anonymous):

y = 6 ?

OpenStudy (kamille):

no

OpenStudy (kamille):

2x + 2y = 6 2y=6-2x y=3-x Now insert everywhere 3-x where is "y" here: x + 3y = −1 and you will find value of x

OpenStudy (anonymous):

okay i think its 2

OpenStudy (kamille):

you are right

OpenStudy (anonymous):

thank u !